OpenStudy (anonymous):

Jack usually mows his lawn in 7 hours. Marilyn can now the same yard in 6 hours. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

solve \[(\frac{1}{7}+\frac{1}{6})t=1\] \[(\frac{6+7}{6\times 7})t=1\] \[t=\frac{6\times 7}{6+7}\]